OpenStudy (anonymous):

Please help will medal and fan!!!!!! A costume designer made 10 costumes that each used 5.8 yards of fabric. If she had 238.2 yards of fabric to start with, about how much fabric does the costume designer have left? Write an equation and estimate the amount of fabric left.

OpenStudy (anonymous):

10 x 5.8 = 58 yards used 238.2 - 58 = ? http://lmgtfy.com/?q=238.2+-+58

OpenStudy (anonymous):

Would that be the equation

OpenStudy (anonymous):

oops, equation: x = 238.2 - 10 x 5.8 x is yards left

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

So what would be the answer??

OpenStudy (anonymous):

x = 238.2 - 10 x 5.8